If |P| = 10, |Q| = 12, and |P ∪Q| = 15, what is |P ∩Q|?pls help with a little explanation.thanks in advance and more power.
Answer by ikleyn(52778)

.
If |P| = 10, |Q| = 12, and |P ∪Q| = 15, what is |P ∩Q|?pls help with a little explanation.thanks in advance and more power.
--------------------------------------------------------------------

If  P  and  Q  are non-intersecting sets of elements,  then it would be  10 + 12 = 22
elements in their union.  But there are only  15  elements in the union.

It means that  P  and  Q  have a non-empty intersection.

Let  x  be the number of elements in the intersection  |P ∩Q|.

Let  p  be the number of elements of  P  that are NOT in  Q,  and
let  q  be the number of elements of  Q  that are NOT in  P.

Then

p + x = 10,           (1)
q + x = 12,           (2)

and

p + q + x = 15       (3)
(for the union  |P ∪Q|).

Add (1) and (2).  You will get

p + q + 2x = 22.       (4).

Now distract (3) from (4).  You will get

x = 22 - 15 = 7.

Answer.  There are  7  elements in the intersection of  P  and  Q:  |P ∩Q| = 7.